OpenStudy (anonymous):

Which line is parallel to the line that passes through the points (-1, 2) and (5, -4)?

OpenStudy (mandre):

Multiple choice?

OpenStudy (mandre):

Do you know how to calculate the slope of the line going through 2 points?

OpenStudy (texaschic101):

slope formula is : slope(m) = (y2 - y1) / (x2 - x1) (-1,2) x1 = -1 and y1 = 2 (5,-4) x2 = 5 and y2 = -4 now lets sub our info into the formula and find the slope slope(m) = (-4 - 2) / (5 - (-1) m = -6 / (5 + 1) m = -6/6= -1 <-- this is the slope of the equation. And since we are looking for a parallel line, the slope will stay the same. Now we will use this formula... y - y1 = m(x - x1) slope(m) = - 1 you can use either of your sets of points, I will use (-1,2) now lets sub it into the formula... y - 2 = -1(x - (-1) y - 2 = -1(x + 1) y - 2 = - x - 1 y = -x - 1 + 2 y = -x + 1 <-- your line